OpenStudy (anonymous):

why does x^2+y^2 = 4-c have no real solutions? (book example about functions of several variables and graphing functions)

OpenStudy (anonymous):

\[\forall c \le 4, x^{2}+y ^{2} = 4 - c, is the equation for a circle with radius = \sqrt{4-c}\]